OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Find the slope which'll be y2-y1/ x2-x1 = 1 Then plug either point into the equation y-y1=m(x-x1) Where m is your slope. So y--4=1(x-0) y+4=1x-0 Add 4 To the other side and there ya go.

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

y=x-4??

OpenStudy (anonymous):

thank you both